# Physics-Informed Neural Networks for Enhanced Interface Preservation in Lattice Boltzmann Multiphase Simulations

### v1

@@ I Introduction
Interface preservation is particularly crucial in applications such as microfluidic device design, where droplet generation and manipulation depend on accurate representation of surface tension effects; in materials science, where phase boundaries dictate material properties; and in geological simulations, where immiscible flow through porous media determines oil recovery efficiency. In each of these scenarios, numerical diffusion can lead to non-physical behavior, inaccurate predictions of breakup or coalescence events, and incorrect estimates of interfacial forces [6, 7].
@@ IV-B Comparative Performance
IV-C Visual Comparison
The most comprehensive comparison is presented in Figure 4, which shows side-by-side views of pure LBM and PINN-LBM results. The superior interface preservation capabilities of the PINN-LBM approach are clearly visible in the gradient magnitude field (right column), where the interface region appears more continuous, uniform, and distinct in the PINN-LBM case compared to the more diffuse and irregular boundary in the pure LBM simulation.
Beyond the interface quality, the velocity fields (middle column) reveal important differences in the flow dynamics. The pure LBM velocity field shows stronger spurious currents near the interface - an undesirable numerical artifact common in multiphase LBM that can lead to non-physical behavior and eventual breakdown of the simulation. In contrast, the PINN-LBM velocity field displays a smoother, more physically consistent pattern with reduced spurious currents, demonstrating that the neural network not only preserves interface sharpness but also improves the overall physical accuracy of the simulation.
@@ V Conclusions
Future work will explore extending this approach to three-dimensional simulations, adapting the PINN architecture for more complex interfacial phenomena, and optimizing the computational efficiency of the coupled system. We also intend to investigate the application of transfer learning techniques to reduce the training overhead for similar simulations, and to explore adaptive coupling strategies that dynamically adjust the frequency of neural network intervention based on local interface properties [22, 17, 23].
VI Acknowledgments
The authors gratefully acknowledge the support from the computational resources provided for this research.

### v2

@@ I Introduction
Interface preservation is particularly crucial in applications such as microfluidic device design, where droplet generation and manipulation depend on accurate representation of surface tension effects; in materials science, where phase boundaries dictate material properties; and in geological simulations, where immiscible flow through porous media determines oil recovery efficiency. In each of these scenarios, numerical diffusion can lead to non-physical behavior, inaccurate predictions of breakup or coalescence events, and incorrect estimates of interfacial forces [6, 7]. Similar challenges have been observed in pharmaceutical applications, where accurate modeling of multiphase flows is essential for predicting protein stability during syringe injection [8] and tablet disintegration processes [9, 10].
@@ IV-B Comparative Performance
IV-C Visual Analysis and Physical Interpretation
The most comprehensive comparison of the two methods is presented in Figure 4, which provides visual evidence of the improvements achieved by the PINN-LBM approach. The PINN-LBM method maintains a more uniform density distribution within each phase, with fewer artifacts near the interface. The velocity magnitude shows smoother transitions across the interface, indicating better physical consistency. The interface region appears more continuous and well-defined in the PINN-LBM case, with reduced numerical diffusion.
The quantitative improvements observed in the metrics are directly reflected in the visual results. The PINN-LBM approach not only maintains sharper interfaces but also preserves the physical characteristics of the flow more accurately. This is particularly important for applications where interface dynamics play a crucial role, such as:
  • Microfluidic Devices: Where precise control of droplet formation and manipulation is required.
  • Materials Processing: Where phase boundaries determine material properties and performance.
  • Geological Simulations: Where accurate representation of multiphase flow through porous media is essential. Recent work in this area has demonstrated the importance of interface preservation for modeling fracturing in coal treated by liquid nitrogen freeze-thaw cycles [27, 28].
@@ V Conclusions
V-A Future Directions
Building on the promising results of our PINN-LBM approach, we have identified several exciting directions for future research that will further enhance the capabilities and applicability of this framework:
  • Advanced Multiphysics Integration: Future developments will incorporate additional physical phenomena such as heat transfer, phase change, and chemical reactions into the PINN-LBM framework. This multiphysics integration will enable more comprehensive simulations of complex systems, similar to recent advancements in boundary control for partial differential equations [31, 32].
  • Adaptive Training Strategies: We plan to develop adaptive training strategies that focus computational resources on regions with significant interface dynamics. This approach will include spatially varying the weighting of different loss components and implementing local refinement of the neural network representation. Such adaptive optimization techniques have shown promise in large-scale neural network training [33] and could be adapted for our PINN-LBM framework.
  • Multi-component Systems: Future work will extend the approach to systems with more than two phases, enabling simulation of complex phenomena such as emulsions, foams, and multiphase flows with surfactants. This will require developing new formulations of the interface loss term that can handle multiple interacting interfaces.